OpenStudy (anonymous):

PLEASE HELP! At the given point fine the line that is normal to the curve. Y^4+x^3=y^2+10x, normal at (0,1). I have the answer but I need the steps!

OpenStudy (anonymous):

find*

zepdrix (zepdrix):

Still stuck on this one or no?

OpenStudy (anonymous):

No, I got this one!

OpenStudy (anonymous):

Wait I need this one, sorry!

zepdrix (zepdrix):

lol

OpenStudy (anonymous):

I thought it was the same one haha

zepdrix (zepdrix):

\[\large y^4+x^3=y^2+10x\] \[\large y'(x,y)=m\] M gives us the slope of the line `tangent` to our function. Do you know the relationship between the `tangent` and `normal` lines?

OpenStudy (anonymous):

Sort of, like its a perpendicular line? Or am I totally off?

zepdrix (zepdrix):

Good, yes :) their perpendicular to one another. To get the slope of the normal line, we'll take the `negative reciprocal` of our tangent slope. So let's start by finding \(\large y'\).

OpenStudy (anonymous):

Okay, so where I was stuck on was the powers, how do I get it in y equals form since there are like 3 different powers used

zepdrix (zepdrix):

Each time we differentiate a y term, a y' will pop out right? So you stuck at this point? \[\large 4y^3y'+3x^2=2yy'+10\]

OpenStudy (anonymous):

yes!

zepdrix (zepdrix):

We want to solve for y'. Let's start by getting all the (y')'s to one side. We'll subtract 2yy' from each side, and subtract 3x^2 from each side,\[\large 4y^3y'-2yy'=10-3x^2\] From here we can `factor` y' out of each term on the left,\[\large y'(4y^3-2y)=10-3x^2\]Then to solve for y', we divide both sides by that big mess,

OpenStudy (anonymous):

Okay, so I understand up to here

OpenStudy (anonymous):

That would be y'=10-3x^2/4y^3-2y

zepdrix (zepdrix):

Ok good :) Since y' is a function of `both` x and y, let's write it like this, \[\large y'(x,y)=\frac{10-3x^2}{4y^3-2y}\] The slope of the line tangent to our function at a particular point will be given by,\[\large y'(0,1)\] So plug that in and see what you get.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

So the equation I get is y=5x+1?

zepdrix (zepdrix):

So if we were asked to find an equation for the line `tangent to the curve`, we would say, "ok so 5 is the slope of our line" And we would write it in this form,\[\large y=mx+b\] But in this problem we want an equation for the line `normal to the curve`. So we'll need to take the negative reciprocal of the slope value we found. That will be the M in our normal line equation.

OpenStudy (anonymous):

So the negative reciprocal is -1/5? So our final answer is y=-1/5x+1?

zepdrix (zepdrix):

You were able to find the y-intercept with no problem? Ok cool. Yah I think that's right :) Lemme check on wolfram real quick to make sure I didn't make a silly mistake somewhere.

zepdrix (zepdrix):

Blah the function is too complicated to input XD lol I'm pretty sure we did that right anyway :D

zepdrix (zepdrix):

Oh do you have an answer key for this one? :3

OpenStudy (anonymous):

No if you say that's right then it is, cause I have the correct answer! I just didn't know the process but I followed what you did and I came out with the same answer! You are so helpful! :) Do you think you could help me on a couple more?

zepdrix (zepdrix):

Turned out correct? Ah good! Sure.

OpenStudy (anonymous):

You are a lifesaver!

zepdrix (zepdrix):

The best way would be to umm, go to your questions and type @zepdrix in the description somewhere. It sends me a notification, making it easier to find the problem.

OpenStudy (anonymous):

Okay I'll do that! so I just type it in the ask a question box?

zepdrix (zepdrix):

Yes that's fine. Or you can Ask the question, and then type it in the posts area.